OpenStudy (anonymous):

how to find the LCM of 5, 6, 15

hero (hero):

30

OpenStudy (anonymous):

15 = 3*5 6 = 3*2 5 lcm = 3*5*2 - pick out only one of the duplicates(triplicates etc) and lone numbers then multiply them